OpenStudy (fabgirl1994):

Given the following functions f(x) and g(x), solve (f + g)(5) and select the correct answer below: (1 point) f(x) = 8x − 13 g(x) = x − 9 a 23 b 27 c 31 d 40

OpenStudy (maddy1251):

Are you good with these or no?

OpenStudy (maddy1251):

@fabgirl1994

OpenStudy (fabgirl1994):

I'm not sure how to do them. My lesson will not load.

OpenStudy (maddy1251):

Okay! It's been a little while since I have done these but I can help you. Just disregard any errors that I may possibly have. xD

OpenStudy (maddy1251):

When we have \[f(x)= x ; g(x)=x\] and we want to combine them ; \[f(g(x))\]

OpenStudy (maddy1251):

or in this case we are adding them together so we have f(x) = 8x − 13 g(x) = x − 9

OpenStudy (fabgirl1994):

Right

OpenStudy (maddy1251):

Okay, you are following along. :) This is rather simple since are just adding the two functions together.

OpenStudy (fabgirl1994):

Yes

OpenStudy (maddy1251):

f(x) = 8x − 13 g(x) = x − 9 (f+g)= (8x-13 + x - 9 )

OpenStudy (maddy1251):

Sorry for taking so long, OS isn't letting me copy the equations to make this go by faster

OpenStudy (maddy1251):

Can you add together like terms for me?

OpenStudy (fabgirl1994):

You're fine. I'm following along.

OpenStudy (fabgirl1994):

Would it be: 8x-3?

OpenStudy (fabgirl1994):

I got that because 8x and and x combined. Then I did 13-9=3

OpenStudy (maddy1251):

(f+g)= (8x-13 + x - 9 ) Let me see :) (8x-13+ 1x - 9 (9x - 22) is what I get.

OpenStudy (maddy1251):

@mathmale am I doing everything correctly?

OpenStudy (mathmale):

@Maddy1251 : Your f(g(x)) does not describe the operation required in this problem. Please explain what the goal of the problem is, if it's not f(g(x)).

OpenStudy (fabgirl1994):

OKay, i get that. I wasn't sure if I added the 1x on or not to the 8. And I seen a + and a - after the 13 so wasn't sure if I added or subtracted 13 and 9.

OpenStudy (maddy1251):

@mathmale I understood after typing f(g(x)) was unrelated to this. This is adding two functions together.

OpenStudy (fabgirl1994):

So how do I get an answer from 9x-22?

OpenStudy (mathmale):

Yes, that's right; it's addition.

OpenStudy (maddy1251):

@fabgirl1994 you have x = 5. You substitute.

OpenStudy (maddy1251):

@mathmale yes you do. It my all my fault mentioning that. I was thinking one concept when the concept didn't need applied.

OpenStudy (mathmale):

Right. Add functions f and g. THEN, let x=5.

OpenStudy (maddy1251):

well... (f+g)(x) where ; x=5

OpenStudy (fabgirl1994):

So x = 5, but thats not an answer choice, what now?

OpenStudy (maddy1251):

Yes, that is what I had her do. 9x - 22.

OpenStudy (fabgirl1994):

I have the answer choices 23, 27, 31 and 40.

OpenStudy (maddy1251):

(f+g)(5) = 9x-22 = ____ ; when x=5

OpenStudy (maddy1251):

9(5)-22.

OpenStudy (fabgirl1994):

Thanks! I have a few more, Ill post the question.

OpenStudy (mathmale):

Yes, I agree with that. You have successfully combined the 2 functions f and g. Fabgirl1994: Let's concentrate on understanding, not just on the answer. (f+g)(x)=9x-22. what happens if you now substitute x=5 into the right side of this equation?

OpenStudy (fabgirl1994):

I got 23 :)

OpenStudy (maddy1251):

There ya go fabgirl.

OpenStudy (maddy1251):

Thanks for catching my fault before @mathmale

OpenStudy (mathmale):

Mind posting further questions as fresh, new posts? The current one has been around for a while. ;)